# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v22

@@ 3.23 Sinhala Sign Language
Meyler [555] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language. Further, [556] claim that SSL was derived from British Sign Language (BSL). However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL. In this section, we will discuss works that have used either term. But when relevant we will point out the works that claimed to work on Sri Lankan Sign Language as opposed to Sinhala Sign Language. As far as we can observe, there does not seem to be a technical or functional difference between the two.
@@
In the domain of Sinhala sign language, Liyanaarachchi et al. [557] claims to have created a signing dataset for the Sinhala sign language however only the abstract of their work can be publicly accessible. DISSANAYAKE et al. [558] also claims to have created a database of gestures that are included in Sinhala sign language. Charuka et al. [143] have published a dynamic word-level Sinhala Sign Language video dataset of 50 classes. The dataset contains 1110 videos.
@@
Priyankara et al. [559] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [560] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [561], the approach by Dilakshan and Priyadarshana [562], as well as the approach by Peiris [563] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [564] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [565] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [566]. The study by Charuka et al. [143] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [567] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [568] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. The study by Fernando et al. [569] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [570] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [571] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [572] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [573]. The work by Perera [574] uses CNN for Sri Lanka Sign Language recognition. Strides have been made in the domains of computer interpreting for written Sinhala [575] and animation of finger-spelled words and number signs [576]. The work by Idushan et al. [577] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
Kumar et al. [578] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. Both the approaches by Perera and Kumarika [579] and Herath and Ishanka [580] integrated MediaPipe [581] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [579] only used LSTM while Herath and Ishanka [580] used LSTM, CNN, and CNN-LSTM. Rishan et al. [556] uses Leap Motion Controller119119119https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [582] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [593] and later by Smith [594] using an XGB classifier and a CRF model building on their previous work [595], which analysed such data. Shanmugalingam and Sumathipala [596] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [465]. They later used it for NMT in Sinhala-English Code-Mixed text [464]. As discussed in Section 3.10Rathnayake et al. [223] used adapter-based [224, 225, 226, 227, 228, 229] fine-tuning on XLM-R [230], for classifying code-mixed and code-switched Sinhala text. Hettigoda [233] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [253] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [235] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [547] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [404] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [597] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by uda [598] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection.
@@ 3.27 Language Acquisition and Evaluation
3.28 Embeddings and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [604]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [101] has published two Sinhala BERT models: SinBERT-large120120120https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small121121121https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo122122122https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [605] on the OSCAR [144, 145] dataset.
Üstün et al. [606] introduced Aya, a massively multilingual generative language model123123123https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [148].
Wickramasinghe and de Silva [138] used Sinhala-English dictionary data [137] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub124124124https://bit.ly/3t3SKu7.
@@ 3.29 Miscellaneous Applications
In this section, we discuss NLP tools and research which are either hard to categorize under above sections or reasonably involve multiples of them.
Dissanayake and Hettige [607] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [608] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha. On the other hand, Amarasinghe and Ranathunga [609] used an ontology-based approch to generate Sinhala essay questions. The work by Liyanage and Ranathunga [610, 611] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages. The follow-up work by Niyarepola et al. [612] used pre-trained mBART [471] and mT5 [299, 613] models for the same task and showed better results. The study by Kao and Ilmini [614] used LSTM to generate Sinhala lyrics. Bandara et al. [615] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala. However, they did not train a GTP-2 model in Sinhala. Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API.
@@
Kumari and Hettiarachchi [626] proposed an algorithm for Sinhala topic modelling based on LDA [232] and RAKE [627]Batawalaarachchi [628] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [629] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [630] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [631] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [632] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [633] have created a small Sinhala treebank of 100 sentences. Burchell et al. [136] show that they obtain better results for the LID task for most of the languages compared to NLLB [127]. However, for Sinhala the results by Burchell et al. [136] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Minixhofer et al. [634] have used Sinhala from OPUS100 [124] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [635] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [613] and mT5 [299, 613] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [636]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [637]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [638]. Senarathne [639] introduced a model to translate Sinhala pseudocode to computer code written in C#. Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [597] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [640] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets.
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [655, 656, 657, 658, 659] or foreigners [660, 193, 661]125125125Note that [661] is an extension of [193].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.

### v23

@@ 3.23 Sinhala Sign Language
Meyler [578] claim that there is no such thing as a Sinhala Sign Language or a Tamil Sign Language and there is only Sri Lankan Sign Language. Further, [579] claim that SSL was derived from British Sign Language (BSL). However, the literature uses Sri Lankan Sign Language and Sinhala Sign Language interchangeably with the same acronym, SSL. In this section, we will discuss works that have used either term. But when relevant we will point out the works that claimed to work on Sri Lankan Sign Language as opposed to Sinhala Sign Language. As far as we can observe, there does not seem to be a technical or functional difference between the two. The work by Chithrani et al. [580] compares a few Sinhala sign language recognition works [581, 582, 583, 584, 585, 586, 579, 587] against work conducted for other languages.
@@
In the domain of Sinhala sign language, Liyanaarachchi et al. [588] claims to have created a signing dataset for the Sinhala sign language however only the abstract of their work can be publicly accessible. DISSANAYAKE et al. [589] also claims to have created a database of gestures that are included in Sinhala sign language. Charuka et al. [147] has published a dynamic word-level Sinhala Sign Language video dataset of 50 classes. The dataset contains 1110 videos.
@@
Priyankara et al. [590] have conducted a short survey on the available Sinhala sign language recognition systems. They look at 12 studies and conclude that image processing with CNN is the most used technique for Sign Language translation. Along with a Sinhala sign language recognition system. Wijegoonaratna [591] has created a neural network-based approach for real-time Sinhala sign language gesture recognition. The approach by Hettiarachchi and Meegama [587], the approach by Dilakshan and Priyadarshana [592], as well as the approach by Peiris [585] have used CNN to recognize the fingerspelling alphabet of Sinhala sign language. Perera and Jayalal [593] also has used CNN to translate Sinhala sign Language to Sinhala text. However, they have also explicitly used Scale Invariant Feature Transform (SIFT) [594] features. A study for the limited use case of translating 15 Sinhala signs to text was conducted by Fernando and Wimalaratne [581]. The study by Charuka et al. [147] introduces a novel skeleton-based sign language recognition method named ProtoSign built upon prototypical networks [595] to do few-shot learning for Sinhala sign language recognition. The study by Haputhanthri et al. [596] used ResNet followed by a transformer encoder with multi-head attention to recognize sentence-level Sinhala sign language. The study by Fernando et al. [597] introduced a mobile app capable of translating Sinhala Sign language to text. They further enhance the results of their system with a facial emotion detection system created with CNNs. The study by Gamage et al. [598] also combines Sinhala Sign language recognition and emotion recognition. Krishnananthan et al. [599] introduced HANDTALK web-based application designed to teach sign-language. As Such it has a component that translates the sign language gestures of the learner into text. In their study Gedaragoda et al. [600] used traditional machine learning approaches to recognize Sinhala Sign Language. A pose-based method for Sinhala Sign Language recognition was introduced by Indatissa [601]. The work by Perera [602] uses CNN for Sri Lanka Sign Language recognition. The work by Nadeesha and Wasalthilaka [603] used classical machine learning techniques to recognize 40 Sinhala signs from 100 pose videos. They reported a success rate of 86.7%. Strides have been made in the domains of computer interpreting for written Sinhala [604] and animation of finger-spelled words and number signs [605]. The work by Idushan et al. [606] claims to be able to convert Sinhala audio to SSL and also recognize SSL and interpret dynamic signs.
@@
The work by Samarasinghe et al. [582] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [607] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [586] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [584] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [608] and Herath and Ishanka [609] integrated MediaPipe [610] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [608] only used LSTM while Herath and Ishanka [609] used LSTM, CNN, and CNN-LSTM. Rishan et al. [579] uses Leap Motion Controller127127127https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [583] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [621] and later by Smith [622] using an XGB classifier and a CRF model building on their previous work [623], which analysed such data. Shanmugalingam and Sumathipala [624] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [485]. They later used it for NMT in Sinhala-English Code-Mixed text [484]. As discussed in Section 3.10Rathnayake et al. [232] used adapter-based [233, 234, 235, 236, 237, 238] fine-tuning on XLM-R [239], for classifying code-mixed and code-switched Sinhala text. Hettigoda [242] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [262] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [244] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [570] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [420] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [625] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by Udawatta et al. [626] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection. Uthpala and Thirukumaran [264] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
@@ 3.27 Language Acquisition and Evaluation
3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [632]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [104] has published two Sinhala BERT models: SinBERT-large128128128https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small129129129https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo130130130https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [633] on the OSCAR [148, 149] dataset.
Üstün et al. [634] introduced Aya, a massively multilingual generative language model131131131https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [152].
Wickramasinghe and de Silva [141] used Sinhala-English dictionary data [140] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub132132132https://bit.ly/3t3SKu7.
Jayakody and Dias [635] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude133133133https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o134134134https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)135135135https://llama.meta.com/docs/get-started/ and Mistral 7B136136136https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2137137137https://falconllm.tii.ae/ and Gemini138138138https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
3.29 Text Simplification
The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [636] included the NSINA [105] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data139139139https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data140140140https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [637] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [638] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [639] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [640] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
@@ 3.30 Miscellaneous Applications
In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them.
Dissanayake and Hettige [641] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [642] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha. Romero et al. [154] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data141141141https://huggingface.co/datasets/afaji/cvqa and leader board142142142https://eval.ai/web/challenges/challenge-page/2305/overview. On the other hand, Amarasinghe and Ranathunga [643] used an ontology-based approach to generate Sinhala essay questions. The study by Kao and Ilmini [644] used LSTM to generate Sinhala lyrics. Bandara et al. [645] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala. However, they did not train a GTP-2 model in Sinhala. Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API. The work by Liyanage and Ranathunga [646, 647] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages. The follow-up work by Niyarepola et al. [648] used pre-trained mBART [491] and mT5 [310, 649] models for the same task and showed better results.
@@
Kumari and Hettiarachchi [660] proposed an algorithm for Sinhala topic modelling based on LDA [241] and RAKE [661]Batawalaarachchi [662] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [663] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [664] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [665] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [666] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [667] have created a small Sinhala treebank of 100 sentences. Burchell et al. [139] show that they obtain better results for the LID task for most of the languages compared to NLLB [130]. However, for Sinhala the results by Burchell et al. [139] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Minixhofer et al. [668] have used Sinhala from OPUS100 [127] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [669] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [649] and mT5 [310, 649] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [670]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [671]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [672]. Senarathne [673] introduced a model to translate Sinhala pseudocode to computer code written in C#. Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [625] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [674] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [675] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. The study by [208] used NER to cluster Sinhala news articles.
@@ 4 Primary Sources
A number of NLP applications have utilized first sources intended to teach children [690, 691, 692, 693, 694] or foreigners [695, 200, 696]143143143Note that [696] is an extension of [200].. This makes sense given that an introduction written for children would start with basic principles and thus be ideal for crafting rule based NLP systems and an introduction written for foreigners would have Sinhala language described in terms of English, making easy the process of rule based translation of English NLP tools to Sinhala.